# Gravitational $D$-Form Factor: The $\sigma$-Meson as a Dilaton confronted with Lattice Data

### v1

@@ Appendix A The Gell-Mann Lévy linear $\sigma$-model – a study case
The NLO form factor can be parameterised as
@@
In the following, we define the model, present the NLO computation, and carry out the analytic continuation. We then use the result to assess the impact of the broad $\sigma$-resonance on the form factor parameterisation.
@@ A.1 The linear $\sigma$-model Lagrangian in the broken phase
The linear $\sigma$-model is a most formidable model, first introduced by Schwinger [107] to implement $SU(2)_{L}\times SU(2)_{R}$ invariance in the strong-interaction sector. It was then refined by Gell-Mann and Lévy [106] to include the mechanism of spontaneous symmetry breaking, which leads to massless pions and mass generation of the nucleon, referred to as the Yukawa-mechanism. The renormalisation was worked out in ref. [108], providing an important precursor to the renormalisation of the electroweak sector of the Standard Model. It should be emphasised that the $\sigma$-particle, in this context, is not to be confused with the $\sigma$-meson of dilaton chiral perturbation theory. In fact, integrating out a potentially heavy $\sigma$, in the linear model, leads to chiral perturbation theory, i.e. the non-linear $\sigma$-model. This allows to determine the low energy constants explicitly. Gasser and Leutwyler [109] make the point that, since the predictions differ from the values observed in nature, the linear $\sigma$-model  cannot be considered a viable theory of the strong interactions.
The Gell-Mann Lévy linear $\sigma$-model with pion mass [106], reads
@@
The renormalisation [108] of the linear $\sigma$-model is involved, but since we do not aim to match it to experiment we may adapt the $\overline{\text{MS}}$-scheme within dimensional regularisation. In fact, we slightly modify it for the self-energies to cancel tadpole diagrams (for $\mu=m_{\sigma}$), so that we may simply ignore them. The concrete values used for (A.9) will be discussed in section A.4.
@@ A.2 Relevant NLO corrections
where $A_{0}=I_{0}$, $B_{0}=I_{1}$ and $C_{0}=I_{2}$. For the momentum routing of the triangle function $C_{0}$ we use the LoopTools conventions [105].
@@ A.2.1 Self-energy corrections $\Sigma(s)$
which contributes a large part the imaginary part of $s_{\sigma}$. Equation (A.15) can be solved numerically, or alternatively approximated perturbatively by substituting $s=m_{\sigma}^{2}$ into the self-energy [^9] This expression coincides with Eq. (1) in [110], up to the choice of scheme and with the $\Sigma_{g}$ term omitted.
@@ A.3 Second-sheet analytic continuation of the form factor
where $\text{disc}[f(s)]=f(s+i0)-f(s-i0)$ ensures continuity across the cut, $f^{(II)}(s-i0)=f(s+i0)$. Concretely, we have the imaginary parts of the Passarino-Veltman functions analytically, while the real parts can be evaluated numerically through the LoopTools package [105] (for the $C_{0}$ function). Fortunately, this proves sufficient since i) the form factors satisfy real analyticity which implies $\text{disc}[f(s)]=2i\textrm{Im}f(s)$ by Schwartz’s reflection principle and ii) a representation for $f(s)$ valid on the physical sheet is given by the standard dispersion relation
@@
The imaginary parts of $B_{0}$ and $C_{0}$ are provided in Itzykson and Zuber [111], although the expression for $C_{0}$ contains significant typos in the Källén function, which we corrected and verified against LoopTools. On the real line we express them as
@@ A.4 Concluding the form-factor study in the linear $\sigma$-model
3 0.27 0.25 -0.036 $0.24e^{0.6^{\circ}i}$ 0.13+0.03i-0.13+0.03i
5 0.33 0.30 -0.059 $28e^{-4.8^{\circ}i} 0.0$7+0.02i-0.07+0.02i
7 0.38 0.34 -0.084 $0.30e^{-9.7^{\circ}i}$ 0.140.10i-0.14-0.10i
9 0.42 0.38 -0.112 $36e^{-16.1^{\circ}i} 0.0$80.25i-0.08-0.25i
11 0.46 0.40 -0.142 $42e^{-19.1^{\circ}i} 0.0$00.34i-0.00-0.34i
13 0.49 0.42 -0.174 $49e^{-20.7^{\circ}i} 0.0$70.40i0.07-0.40i
15 0.53 0.44 -0.208 $0.55e^{-21.6^{\circ}i}$ 0.140.45i0.14-0.45i
@@ Appendix B A Multipole Expansion in Momentum Space
$G_{\varphi}(q^{2})=\frac{1}{q^{2}-m_{\mathrm{E},\varphi}^{2}}\Big{(}r_{\mathrm{E},\varphi}+\sum_{n\geq 1}{\cal E}_{n}\Big{)}\;,\qquad{\cal E}_{n}=\frac{\Omega_{n}}{(q^{2}-m_{\mathrm{E},\varphi}^{2})^{n}}\;,$ (B.3)
with moments defined as
@@
which converge if $\rho_{\varphi}$ has finite support. For sufficiently Euclidean $q^{2}$, the hierarchy $|{\cal E}_{n+1}/{\cal E}_{n}|\ll 1$ is expected to hold, and in practice this may extend to all $q^{2}<0$. Note that $\Omega_{0}=-r_{\mathrm{E},\varphi}$ which explains the change in sign in the main $q^{2}$-denominator. This expansion is analogous to the multipole expansion in classical electrodynamics, though here it takes place in momentum rather than coordinate space. The analogue of the charge-distribution centre is played by $m_{\mathrm{E},\varphi}$, the approximate centre of the $\rho_{\varphi}(s)$-distribution. The table and figure in appendix A.4 show that the assertions made are true in the linear $\sigma$-model for specific parameter ranges. See also the discussion in section 3.1.

### v2

@@ Appendix A The Gell-Mann Lévy linear $\sigma$-model - a Toy Model
The NLO form factor can be parametrised as
@@
In the following, we define the model, present the NLO computation, and carry out the analytic continuation. We then use the result to assess the impact of the broad $\sigma$-resonance on the form factor parametrisation.
@@ A.1 The linear $\sigma$-model Lagrangian in the broken phase
The linear $\sigma$-model is a most formidable model, first introduced by Schwinger Schwinger:1957em to implement $SU(2)_{L}\times SU(2)_{R}$ invariance in the strong-interaction sector. It was then refined by Gell-Mann and Lévy Gell-Mann:1960mvl to include the mechanism of spontaneous symmetry breaking, which leads to massless pions and mass generation of the nucleon, referred to as the Yukawa-mechanism. The renormalisation was worked out in ref. Gervais:1969zz , providing an important precursor to the renormalisation of the electroweak sector of the Standard Model. It should be emphasised that the $\sigma$-particle, in this context, is not to be confused with the $\sigma$-meson of dilaton chiral perturbation theory. In fact, integrating out a potentially heavy $\sigma$, in the linear model, leads to chiral perturbation theory, i.e. the non-linear $\sigma$-model. This allows to determine the low energy constants explicitly. Gasser and Leutwyler Gasser:1983yg make the point that, since the predictions differ from the values observed in nature, the linear $\sigma$-model cannot be considered a viable theory of the strong interactions.
The Gell-Mann Lévy linear $\sigma$-model with pion mass Gell-Mann:1960mvl , reads
@@
The renormalisation Gervais:1969zz of the linear $\sigma$-model is involved, but since we do not aim to match it to experiment we may adapt the $\overline{\text{MS}}$-scheme within dimensional regularisation. In fact, we slightly modify it for the self-energies to cancel tadpole diagrams (for $\mu=m_{\sigma}$), so that we may simply ignore them. The concrete values used for (A.9) will be discussed in section A.4.
@@ A.2 Relevant NLO corrections
where $A_{0}=I_{0}$, $B_{0}=I_{1}$ and $C_{0}=I_{2}$. For the momentum routing of the triangle function $C_{0}$ we use the LoopTools conventions Hahn:1998yk .
@@ A.2.1 Self-energy corrections $\Sigma(s)$
which contributes a large part the imaginary part of $s_{\sigma}$. Equation (A.15) can be solved numerically, or alternatively approximated perturbatively by substituting $s=m_{\sigma}^{2}$ into the self-energy [^11] This expression coincides with Eq. (1) in Masjuan:2008cp , up to the choice of scheme and with the $\Sigma_{g}$ term omitted.
@@ A.3 Second-sheet analytic continuation of the form factor
where $\text{disc}[f(s)]=f(s+i0)-f(s-i0)$ ensures continuity across the cut, $f^{(II)}(s-i0)=f(s+i0)$. Concretely, we have the imaginary parts of the Passarino-Veltman functions analytically, while the real parts can be evaluated numerically through the LoopTools package Hahn:1998yk (for the $C_{0}$ function). Fortunately, this proves sufficient since i) the form factors satisfy real analyticity which implies $\text{disc}[f(s)]=2i\textrm{Im}f(s)$ by Schwartz’s reflection principle and ii) a representation for $f(s)$ valid on the physical sheet is given by the standard dispersion relation
@@
The imaginary parts of $B_{0}$ and $C_{0}$ are provided in Itzykson and Zuber Itzykson:1980rh , although the expression for $C_{0}$ contains significant typos in the Källén function, which we corrected and verified against LoopTools. On the real line we express them as
@@ A.4 Concluding the form-factor study in the linear $\sigma$-model
3 0.27 0.25 0.036-0.036 $0.24e^{0.6^{\circ}i}$ 0.13+0.03i-0.13+0.03i
5 0.33 0.30 0.059-0.059 $28e^{-4.8^{\circ}i} 0.0$7+0.02i-0.07+0.02i
7 0.38 0.34 0.084-0.084 $0.30e^{-9.7^{\circ}i}$ 0.140.10i-0.14-0.10i
9 0.42 0.38 0.112-0.112 $36e^{-16.1^{\circ}i} 0.0$80.25i-0.08-0.25i
11 0.46 0.40 0.142-0.142 $42e^{-19.1^{\circ}i} 0.0$00.34i-0.00-0.34i
13 0.49 0.42 0.174-0.174 $49e^{-20.7^{\circ}i} 0.0$70.40i0.07-0.40i
15 0.53 0.44 0.208-0.208 $0.55e^{-21.6^{\circ}i}$ 0.140.45i0.14-0.45i
@@ Appendix B A Multipole Expansion in Momentum Space
$G_{\varphi}(q^{2})=\int_{0}^{\infty}\frac{ds\,\rho_{\varphi}(s)}{s-q^{2}-i0}=\frac{1}{q^{2}-m_{\mathrm{E},\varphi}^{2}}\Big(r_{\mathrm{E},\varphi}+\sum_{n\geq 1}{\cal E}_{n}\Big)\;,$ (B.3)
where
${\cal E}_{n}=\frac{\Omega_{n}}{(q^{2}-m_{\mathrm{E},\varphi}^{2})^{n}}\;,$ (B.4)
with
@@
converging moments provided $\rho_{\varphi}$ has finite support. For sufficiently Euclidean $q^{2}$, the hierarchy $|{\cal E}_{n+1}/{\cal E}_{n}|\ll 1$ is expected to hold, and in practice this may extend to all $q^{2}<0$. Note that $\Omega_{0}=-r_{\mathrm{E},\varphi}$ which explains the change in sign in the main $q^{2}$-denominator. This expansion is analogous to the multipole expansion in classical electrodynamics, though here it takes place in momentum rather than coordinate space. The analogue of the charge-distribution centre is played by $m_{\mathrm{E},\varphi}$, the approximate centre of the $\rho_{\varphi}(s)$-distribution. The table and figure in appendix A.4 show that the assertions made are true in the linear $\sigma$-model for specific parameter ranges. See also the discussion in section 3.1.
Appendix C Comparison plots
Figure 5 shows the $D$-form factors of the pion (4.6) and the nucleon (4.1) (of the Euclidean pole fits in Fig. 1 with $r_{\text{eff}}=0$), compared to to the $n$-pole fits of the original studies Hackett:2023rif ; Hackett:2023nkr and to the narrow-resonance approximation in refs. Broniowski:2024oyk ; Broniowski:2025ctl .
Figure 5: The $D$-form factor of the nucleon (left) and pion (right) as a function of the momentum transfer $q^{2}$. Comparison between our main parametrisation (green) as also shown in Fig. 1, i.e., Eqs. (4.1) and (4.6) with $r_{\mathrm{eff}}=0$, the $n$-pole fits performed in the original paper by Hackett et al. Hackett:2023rif ; Hackett:2023nkr (blue) and in the meson dominance approach by Broniowski and Ruiz Arriola refs. Broniowski:2024oyk ; Broniowski:2025ctl (red). The fits are compared to the data of refs. Hackett:2023rif ; Hackett:2023nkr .
In both cases agreement at the $1\sigma$-level is found, between the three fits, in the entire data-range. However, we wish to emphasise that the aim of this work is not to obtain the best description of this data, but the physical interpretation in terms of dilaton effective field theory (as discussed in section 2).
Next we consider the trace of the energy-momentum tensor $\Theta^{N,\pi}(q^{2})$ which is obtained by taking the trace of Eq. 2.1 which are normalised as $\Theta^{N,\pi}(0)=2m_{N,\pi}^{2}$ (for $m_{\sigma}\neq 0$).
[…]
In Fig. 6 we show the fitted $\Theta$-form factors for the nucleon and the pion, from the linear combination of the independently fitted $A$-, $J$- and $D$-form factors. [^12] Following refs. Hackett:2023nkr ; Hackett:2023rif , the $A$- and $J$-form factors are fitted using $n$-pole parametrisations, $F(q^{2})=\alpha/(1-q^{2}/\Lambda^{2})^{n}$. Specifically, a monopole parametrisation for the pion and a dipole parametrisation for the nucleon. The nucleon dipole fit-parameters are $\alpha_{A}=0.99(4)$, $\Lambda_{A}=1.37(3)$, $\alpha_{J}=0.50(3)$, $\Lambda_{J}=-1.51(7)$ with correlations: $\rho_{\alpha_{A},\Lambda_{A}}=-0.46$, $\rho_{\alpha_{D},\Lambda_{D}}=0.83$. The pion monopole fit-parameters are $\alpha_{A}=1.00(2)$, $\Lambda_{A}=1.18(2)$ wtith correlations: $\rho_{\alpha_{A},\Lambda_{A}}=-0.14$. Hence, their uncertainties are added in quadrature and equally so for the MIT lattice data. To evaluate the $\Theta$-form factor we take $m_{N}=903\,\mbox{MeV}$ private for the MIT data at $m_{\pi}=170\,\mbox{MeV}$ and $m_{N}=1096\,\mbox{MeV}$ (provided with the data) Wang:2024lrm for the $m_{\pi}=235\,\mbox{MeV}$ $\chi\mathrm{QCD}$-data Wang:2024lrm . Note that for the $\chi\mathrm{QCD}$-simulation only the field strength part of the trace anomaly has been considered $\frac{\beta}{2g}{G^{2}}$, omitting $m_{q}(1+\gamma_{m})\bar{q}q$. Since the nucleon mass is approximately reproduced for different (pion masses) Wang:2024lrm , it is clear that this is the lion-share and thus comparison is pragmatically possible. We do not compare to the $\chi\mathrm{QCD}$ results for the pion since there the omitted quark contributions are too sizeable.